OpenStudy (anonymous):

anyone good at square root? 1.94 2.29 3.6 4.73 5.34 6.33 12.26 13.85 14.52 15.30 16.6 17.13 14.27 15.67 16.99 17.67 18.99 19.75 20.55 21.5 22.39

OpenStudy (anonymous):

what do you want to do with these numbers?

OpenStudy (anonymous):

you know you can just plug them in to calculator to solve them?(~.~)

OpenStudy (anonymous):

yeah i did and then it ends up with an ongoing number i need to know what two numbers multiply to them or plug into them like< 4> iwhich would be two but some i need two numbers for but i dont know what multiplys for then

OpenStudy (anonymous):

like idk what times what is 39

OpenStudy (anonymous):

so, you plug 39^(1/2)

OpenStudy (anonymous):

or sqrt(39)

OpenStudy (anonymous):

in to your calculator. and just rounded it to two decimal place

OpenStudy (anonymous):

if you wants to be accurate, just wright\[\sqrt{39}\]

OpenStudy (anonymous):

i did sqrt and i got 6.244997998

OpenStudy (anonymous):

there is no other ways you know>.<

OpenStudy (anonymous):

im doing algebra online and its not being very formative in giving me instructions

OpenStudy (anonymous):

..what do you want for your answers?

TheSmartOne (thesmartone):

1. \[\sqrt{47}\times \sqrt{2}\]

TheSmartOne (thesmartone):

2.\[\sqrt{29}\times \sqrt{1}\]

TheSmartOne (thesmartone):

3.\[\sqrt{3}\times \sqrt{2}\]

TheSmartOne (thesmartone):

4.\[\sqrt{73}\times \sqrt{1}\]

TheSmartOne (thesmartone):

Do you get it...

TheSmartOne (thesmartone):

\[\sqrt{xy}=\sqrt{x}\times \sqrt{y}\]

OpenStudy (anonymous):

yeah

TheSmartOne (thesmartone):

so try the rest by yourself and I will see if you did it correct..

TheSmartOne (thesmartone):

I just got warned by a moderator because I am just giving you answers...

OpenStudy (anonymous):

umm...well i kind of erased my answer so i have to redo them LOL but okay

TheSmartOne (thesmartone):

yeah cause I got a warning and i don't want to get suspended. Lol he doesn't now thst I explained it to you in the privatmessage b4 you posted all the questions..

OpenStudy (anonymous):

well now he might cause you just said it out loud lol but for the sqrt for 57 i got 7.549834435 so do i round it? but i need two numbers. and sorry about that my fault..

TheSmartOne (thesmartone):

57 is nothing but 3x19 so using \[\sqrt{xy}=\sqrt{x}\times \sqrt{y}\]

OpenStudy (anonymous):

:s

TheSmartOne (thesmartone):

it becomes \[\sqrt{57}=\sqrt{3\times 19}=\sqrt{3}\times \sqrt{19}\]

OpenStudy (anonymous):

o.o

OpenStudy (anonymous):

so.... how can i eliminate choices or numbers to find what equals the number? if that makes sense

TheSmartOne (thesmartone):

5. \[\sqrt{34}=\sqrt{17\times2}=\sqrt{17}\times \sqrt{2}\]

TheSmartOne (thesmartone):

first see if tehy can be divided by prime numbers such as 2, 3, 5, 7...

OpenStudy (anonymous):

ohhh okay let me try a few

OpenStudy (anonymous):

what if they can be ?

OpenStudy (anonymous):

34 can only go into two but its 17 so umm...OH I GET IT

OpenStudy (anonymous):

2x17=34

TheSmartOne (thesmartone):

lets say p is a prime number then \[\sqrt{p}=\sqrt{p}\times \sqrt{1}\]

TheSmartOne (thesmartone):

exactly

OpenStudy (anonymous):

yeah so i just dive each number by 1,2,3,5,7?

OpenStudy (anonymous):

then times it? like 34/17= 2 so 17x2=34?

TheSmartOne (thesmartone):

not necesarilly by all of them if one fits then you have gotten the two numbers.

TheSmartOne (thesmartone):

yeah but don't forget teh square root sign... :)

OpenStudy (anonymous):

right. okay lets see if i can get these then lol cause my calculator is stupid

OpenStudy (anonymous):

what if they all end in big numbers?

OpenStudy (anonymous):

cause for 24 i got all ongoing ones

OpenStudy (tkhunny):

Spreadsheets are wonderful things, but unless you go to a lot of trouble to get more decimal places, things will always LOOK LIKE they end. 1.94 1.392838827718410000000000000 2.29 1.513274595042160000000000000 3.60 1.897366596101030000000000000 4.73 2.174856317093150000000000000 5.34 2.310844001658270000000000000 6.33 2.515949125081820000000000000 12.26 3.501428280002320000000000000 13.85 3.721558813185680000000000000 14.52 3.810511776651530000000000000 15.30 3.911521443121590000000000000 16.60 4.074309757492670000000000000 17.13 4.138840417314980000000000000 14.27 3.777565353504820000000000000 15.67 3.958535082577900000000000000 16.99 4.121892769104990000000000000 17.67 4.203569911396740000000000000 18.99 4.357751713900180000000000000 19.75 4.444097208657790000000000000 20.55 4.533210782657250000000000000 21.50 4.636809247747850000000000000 22.39 4.731807265728390000000000000 The zeros on the end, though, are really fake. It's just as far as the calculator or spreadsheet will go without any special effort.

OpenStudy (anonymous):

im supposed to have two numbers that multiply to the number and nothing goes into 26

TheSmartOne (thesmartone):

26=13x2

TheSmartOne (thesmartone):

also 24=2x12 or 4x6 or 8x3

OpenStudy (anonymous):

oh crap i divided

OpenStudy (anonymous):

lol i didnt multiply urge i forgot the steps can you explain them again

OpenStudy (anonymous):

as like how i would use the number 1,2,3,5,7 would it be like im confusing myself

TheSmartOne (thesmartone):

\[\sqrt{xy}=\sqrt{x}\times \sqrt{y}\]

OpenStudy (anonymous):

lol

TheSmartOne (thesmartone):

so for example\[\sqrt{625}=\sqrt{25}\times \sqrt{25}=5\times5=25\]

OpenStudy (anonymous):

i mean like how i said for 34 i would do 2x17 nvm yeah thats what im trying to say ^

TheSmartOne (thesmartone):

\[\sqrt{34}=\sqrt{17\times2}=\sqrt{17}\times \sqrt{2}\]

OpenStudy (anonymous):

can you explain how i would find it for 26?

OpenStudy (anonymous):

cause i multiplied or divided it by those numbers you gave me and its just a big number